The sum of Tony's age and his older brother's age is 29. The difference between their ages is 5 years. How old is Tony
Answer by rheakod(2)

Ok first identify variables
x=Tony
y= Tony's brother
----------------------------------------------------------------------------------------------------------------
Plug these into the equations
x+y=29
y-x=5
-------------------------------------------------------------------------------------------------------------------
Eliminate to get one variable
x+y=29
-x+y=5
So you would get 2y=34---> y=17
-------------------------------------------------------------------------------------------------------------------
If Tony's brother is 17, then Tony is 5 years younger --->17-5= 12

Tony is 12 years old